OpenStudy (anonymous):

how do you find intergal of 1/(y^2 -1)

OpenStudy (raden):

partial fraction 1/(y^2-1) = A/(y+1) + B/(y-1)

OpenStudy (raden):

u have to find the values of A and B, first

OpenStudy (anonymous):

or trig substitution

OpenStudy (anonymous):

no sorry that would be a terrible idea

OpenStudy (raden):

got it ?

OpenStudy (raden):

1/(y^2-1) = A/(y+1) + B/(y-1) 1/(y^2-1) = ( A(y-1) + B(y+1) )/((y+1)(y-1)) 1/(y^2-1) = ((A+B)y + B-A ) / (y^2-1) A+B = 0 B-A = 1 by using elimination or subtitution, we get B=1/2 and A = -1/2 see ur integral becomes int (-1/2)/(y+1) dy + int(1/2)/(y-1) dy

OpenStudy (raden):

use integral by u-sub

OpenStudy (raden):

for int (-1/2)/(y+1) let u = y+1 du = dy it can be : int (-1/2)/u du = -1/2 ln(u)

OpenStudy (raden):

subtitute back that u = y+1 now, we have int (-1/2)/u du = -1/2 ln(u) = -1/2 ln(y+1)

OpenStudy (raden):

by same idea to solving int(1/2)/(y-1) dy u = y-1 du = du so, we have int(1/2)/u = 1/2 ln(u) = 1/2 ln(y-1)

OpenStudy (raden):

finally, we have the answer : -1/2 ln(y+1) + 1/2 ln(y-1) + c

OpenStudy (raden):

does that make sense ?

OpenStudy (anonymous):

perfect sense ty

OpenStudy (raden):

you're welcome